A parent in a candy store tells her child that she can select 2 pieces of candy from the bins in the store. There are 50 bins with different types of candy. How many different candy choices could the child make if each type of candy is different.
Answer by fractalier(6550)

50 chosen two at a time = 50! / (50-2)!(2!) = 50*49 / 2*1 = 1225 ways